OpenStudy (anonymous):

Find TV if T(9, -2) and V(-5, -6).

OpenStudy (aroub):

midpoint or distance??

OpenStudy (anonymous):

tv=\[\sqrt{452}\]

OpenStudy (aroub):

how did u know if find the midpoint or distance??

OpenStudy (anonymous):

read question carefully..it says to find distance..there is no description abt midpoint

OpenStudy (aroub):

and no description abt distance

OpenStudy (anonymous):

man..TV means itself distance between two points T and V

OpenStudy (anonymous):

how did you get 452 ?

OpenStudy (anonymous):

TV=\[\sqrt{(9+5)^2+(-2+6)^2} =\sqrt{452}\]

OpenStudy (anonymous):

TV could be multiplication of 2 vectors